### v2

# On higher-dimensional symmetric designs

Abstract.
We study two kinds of generalizations of symmetric block designs to higher dimensions, the so-called $\mathcal{C}$-cubes and $\mathcal{P}$-cubes. For small parameters, all examples up to equivalence are determined by computer calculations. Known properties of automorphisms of symmetric designs are extended to autotopies of $\mathcal{P}$-cubes, while counterexamples are found for $\mathcal{C}$-cubes. An algorithm for the classification of $\mathcal{P}$-cubes with prescribed autotopy groups is developed and used to construct more examples. A bound on the dimension of difference sets for $\mathcal{P}$-cubes is proved and shown to be tight in elementary abelian groups. The construction is generalized to arbitrary groups by introducing regular sets of (anti)automorphisms.
Key words and phrases:
higher-dimensional design, symmetric design, difference set
1991 Mathematics Subject Classification:
05B05, 05B10, 05B20
This work has been supported by the Croatian Science Foundation under the project $9752$.
1. Introduction
Two different generalizations of symmetric block designs to higher dimensions have recently been studied in [21] and [22]. A symmetric $(v,k,\lambda)$ design can be represented by its incidence matrix, i.e. by a $v\times v$ matrix $A$ with $\{0,1\}$-entries satisfying
$AA^{t}=(k-\lambda)I+\lambda J.$ (1)
Here, $I$ is the identity matrix and $J$ is the all-ones matrix. An $n$-cube of $(v,k,\lambda)$ designs [21] is an $n$-dimensional $v\times\cdots\times v$ matrix such that its every $2$-section is an incidence matrix of a $(v,k,\lambda)$ design. Sections of dimension $2$ are submatrices obtained by fixing all but two coordinates. This generalization is a special case of W. de Launey’s proper $n$-dimensional transposable designs; see [5, Definitions 2.1, 2.6, 2.7, and Example 2.2]. The set of all $n$-cubes of $(v,k,\lambda)$ designs was denoted by $\mathcal{C}^{n}(v,k,\lambda)$ in [21]. We shall refer to them as $\mathcal{C}$-cubes.
The second generalization was introduced in [22] under the name $(v,k,\lambda)$ projection $n$-cubes. These are $n$-dimensional $v\times\cdots\times v$ matrices with $\{0,1\}$-entries such that every $2$-projection is an incidence matrix of a $(v,k,\lambda)$ design. If $C$ is an $n$-dimensional matrix and $1\leq x<y\leq n$, the projection $\Pi_{xy}(C)$ is the $2$-dimensional matrix with $(i_{x},i_{y})$-entry
$\sum_{1\leq i_{1},\ldots,i_{x-1},i_{x+1},\ldots,i_{y-1},i_{y+1},\ldots,i_{n}\leq v}C(i_{1},\ldots,i_{n}).$
The sum is taken over all $n$-tuples $(i_{1},\ldots,i_{n})\in\{1,\ldots,v\}^{n}$ with fixed coordinates $i_{x}$ and $i_{y}$ in a field of characteristic 00. This definition was inspired by Room squares [10], which are generalized to $n$-dimensional Room cubes in an analogous way. In [22], the set of all $(v,k,\lambda)$ projection $n$-cubes was denoted by $\mathcal{P}^{n}(v,k,\lambda)$. We shall refer to them as $\mathcal{P}$-cubes. Two examples are shown in Figure 1, with three light-sources placed along the coordinate axes so that the projections appear as shadows. These and other pictures in the paper were rendered using the ray tracing software POV-Ray [30].
Figure 1. Cubes from $\mathcal{P}^{3}(7,4,2)$ and $\mathcal{P}^{3}(11,5,2)$.
The purpose of this paper is to study both types of cubes and to compare their properties. For dimension $n=2$, both are just incidence matrices of symmetric designs, but for $n\geq 3$ there are significant differences. For example, it was shown in [22, Theorem 2.9] that the dimension of $(v,k,\lambda)$ projection cubes with $k\geq 2$ is bounded by
$n\leq\frac{vk-1}{k-1}.$ (2)
The largest integer $n$ such that $\mathcal{P}^{n}(v,k,\lambda)$-cubes exists is denoted by $\nu(v,k,\lambda)$. In contrast, the dimension of $\mathcal{C}$-cubes can be arbitrarily large for fixed parameters $(v,k,\lambda)$.
The organization of our paper is as follows. In Section 2, we recall the definitions of isotopy and equivalence of $n$-dimensional incidence cubes. We discuss the numbers of inequivalent cubes in $\mathcal{C}^{n}(v,k,\lambda)$ and $\mathcal{P}^{n}(v,k,\lambda)$ for $k=1$ and $k=2$. The main result in this section is Theorem 2.4, giving a complete enumeration up to equivalence of $\mathcal{P}$-cubes for parameters $(7,3,1)$, $(7,4,2)$, and all dimensions $n$. The proof is a computer calculation based on an algorithm that successively increases the dimension. The largest possible dimensions are $\nu(7,3,1)=7$ and $\nu(7,4,2)=9$.
Autotopies of $\mathcal{C}$- and $\mathcal{P}$-cubes are studied in Section 3. Results about the action of automorphisms of symmetric designs on the points and blocks carry over to the action of autotopies of $\mathcal{P}$-cubes on each coordinate. These results do not hold for autotopies of $\mathcal{C}$-cubes. The main computational result in this section is Theorem 3.8, giving a complete classification of $\mathcal{P}^{n}(11,5,2)$-cubes with nontrivial autotopies. Cubes in $\mathcal{P}^{3}(16,6,2)$ with an autotopy of order $8$ acting semiregularly are classified in Proposition 3.9. Among them are examples with three non-isomorphic $(16,6,2)$ designs as projections, answering a question posed in [22].
In Section 4, we study $\mathcal{P}^{n}(v,k,\lambda)$-cubes constructed from higher-dimensional difference sets. Theorem 4.1 gives a sharper bound than (2) on the dimension: $n\leq v$. Building on results from [22], we compute new values of $\mu_{G}(v,k,\lambda)$, the largest dimension of $(v,k,\lambda)$ difference sets in the group $G$. For elementary abelian groups $G$, Theorem 4.4 shows that the bound is tight, i.e. $\mu_{G}(v,k,\lambda)=v$ holds whenever difference sets exist. Theorem 4.7 generalizes the construction to arbitrary groups $G$ based on the notion of regular sets of (anti)automorphisms. A nice consequence is Corollary 4.8, showing that cyclic $(v,k,\lambda)$ difference sets extend at least to dimension $p$, where $p$ is the smallest prime divisor of $v$.
In the final Section 5, we put forward some observations based on data in Table 4. The table contains numbers of inequivalent $\mathcal{P}^{n}(v,k,\lambda)$-cubes constructed from difference sets. We have only partial explanations for apparent symmetries of the numbers and hope that the observations could lead to new theorems.
Some of our results have traditional formal proofs, while others are proved by computer calculations. There are numerous connections between the two types of results in both directions. In Section 3, formal results about autotopies enable computer classifications of $\mathcal{P}$-cubes with prescribed autotopy groups. In Section 4, computer classifications of higher-dimensional difference sets provide impetus for formal results, notably Theorems 4.4 and 4.7. Higher-dimensional incidence cubes are most easily explored by studying examples and performing experiments on a computer. Tools for examining $\mathcal{C}$- and $\mathcal{P}$-cubes of symmetric designs are available in the GAP [12] package Prescribed automorphism groups [20]. Plenty of examples are given in this paper, as well as in [21, 22].
2. Equivalence and classification
Formally, incidence matrices of symmetric $(v,k,\lambda)$ designs are functions $A:\{1,\ldots,v\}\times\{1,\ldots,v\}\to\{0,1\}$ satisfying equation (1). An important consequence of the equation is non-singularity.
Lemma 2.1.
Incidence matrices of symmetric $(v,k,\lambda)$ designs are invertible.
The lemma implies that the transposed matrix $A^{t}$ is also an incidence matrix of a $(v,k,\lambda)$ design, called the dual design. We refer to the monographs [14, 23] and the book [2] for proofs. Symmetric designs with incidence matrices $A$ and $A^{\prime}$ are isomorphic if there are permutations $\pi_{1},\pi_{2}\in S_{v}$ such that $A^{\prime}(i,j)=A(\pi_{1}^{-1}(i),\pi_{2}^{-1}(j))$, $\forall i,j\in\{1,\ldots,v\}$. This can be written in matrix form as $A^{\prime}=P_{1}AP_{2}^{t}$, where $P_{1}$ and $P_{2}$ are permutation matrices corresponding to $\pi_{1}$ and $\pi_{2}$. We call $A$ and $A^{\prime}$ equivalent if $A^{\prime}$ is isomorphic to $A$ or $A^{t}$. The equivalence classes are orbits of the wreath product $S_{v}\wr S_{2}$ acting on the set of all incidence matrices and represent symmetric designs up to isomorphism and duality.
Both $\mathcal{C}$- and $\mathcal{P}$-cubes are defined as $n$-dimensional incidence matrices of order $v$, i.e. as functions $C:\{1,\ldots,v\}^{n}\to\{0,1\}$ with the Cartesian $n$-ary power of $\{1,\ldots,v\}$ as domain. Isomorphism of $n$-dimensional matrices is called isotopy: $C$ and $C^{\prime}$ are isotopic if there are permutations $\pi_{1},\dots,\pi_{n}\in S_{v}$ such that $C^{\prime}(i_{1},\ldots,i_{n})=C(\pi_{1}^{-1}(i_{1}),\ldots,\pi_{n}^{-1}(i_{n}))$, $\forall i_{1},\ldots,i_{n}\in\{1,\ldots,v\}$. Now the order of the coordinates can be permuted by any $\gamma\in S_{n}$. This is called conjugation:
$C^{\gamma}(i_{1},\ldots,i_{n})=C(i_{\gamma^{-1}(1)},\ldots,i_{\gamma^{-1}(n)}),\kern 5.69054pt\forall i_{1},\ldots,i_{n}\in\{1,\ldots,v\}.$
The cubes $C$ and $C^{\prime}$ are equivalent or paratopic if $C^{\prime}$ is isotopic to a conjugate $C^{\gamma}$. The equivalence classes are orbits of the wreath product $S_{v}\wr S_{n}$ acting on $\mathcal{C}^{n}(v,k,\lambda)$ or $\mathcal{P}^{n}(v,k,\lambda)$. The terminology is borrowed from latin squares [17], where equivalence classes of paratopy are usually called main classes.
The classification problem is to determine the equivalence classes of $\mathcal{C}$- and $\mathcal{P}$-cubes for given parameters $(v,k,\lambda)$ and dimension $n$. We start with the degenerate case $k=1$. Incidence matrices of symmetric $(v,1,0)$ designs are permutation matrices of order $v$. They are all equivalent, since the rows and columns can be permuted to get the identity matrix $I$. This extends straightforwardly to $\mathcal{P}^{n}(v,1,0)$-cubes, which are all equivalent to
$C(i_{1},\ldots,i_{n})=\left\{\begin{array}[]{ll}1,&\mbox{ if }i_{1}=\ldots=i_{n},\\[2.84526pt] 0,&\mbox{otherwise.}\\ \end{array}\right.$
On the other hand, the classification of $\mathcal{C}^{n}(v,1,0)$-cubes is a difficult problem. For $n=3$, they are in $1$-to-$1$ correspondence with latin squares $L=(\ell_{i_{1}i_{2}})$ of order $v$ by $C(i_{1},i_{2},i_{3})=[\ell_{i_{1}i_{2}}=i_{3}]$. The square bracket $[P]$ is the Iverson symbol, taking the value $1$ if $P$ is true, and 00 otherwise [19]. In Figure 2, representatives of the two main classes of latin squares of order $4$ and the corresponding $\mathcal{C}^{3}(4,1,0)$-cubes are shown. The number of main classes of latin squares has been determined by computer calculations up to $v=11$ [13, 25].
$\left(\begin{array}[]{llll}1&2&3&4\\ 2&3&4&1\\ 3&4&1&2\\ 4&1&2&3\end{array}\right)\kern 34.1433pt\left(\begin{array}[]{llll}1&2&3&4\\ 2&1&4&3\\ 3&4&1&2\\ 4&3&2&1\end{array}\right)$
Figure 2. Two $\mathcal{C}^{3}(4,1,0)$-cubes and the corresponding latin squares of order $4$.
Cubes in $\mathcal{C}^{n}(v,1,0)$ of arbitrary dimension $n$ are in $1$-to-$1$ correspondence with latin hypercubes of order $v$ and dimension $n-1$. There have been several definitions of latin hypercubes in the literature; the suitable definition for our purpose is used in [27]. A latin hypercube of order $v$ and dimension $n$ is a $v\times\cdots\times v$ matrix with entries from $\{1,\ldots,v\}$, such that every $1$-section contains every symbol exactly once. The set of all such objects is denoted by $\mathcal{H}^{n}_{v}$. In [27], classification is performed for $n=3$, $v\leq 6$ and $n=4,5$, $v\leq 5$.
For $k=2$, the only feasible parameters of symmetric designs are $(3,2,1)$. The number of $\mathcal{C}^{n}(3,2,1)$ cubes can be determined by complementation, i.e. by exchanging $0\leftrightarrow 1$. This is a bijection between $\mathcal{C}^{n}(v,k,\lambda)$ and $\mathcal{C}^{n}(v,v-k,v-2k+\lambda)$.
Proposition 2.2.
The total number of $\mathcal{C}^{n}(3,2,1)$-cubes is $3\cdot 2^{n-1}$ and they are all isotopic.
Proof.
By complementation, it suffices to count cubes in $\mathcal{C}^{n}(3,1,0)$. They are in $1$-to-$1$ correspondence with latin hypercubes of order $v=3$ and dimension $n-1$. The number of such hypercubes is known to be $|\mathcal{H}_{3}^{n-1}|=3\cdot 2^{n-1}$, see [27, page 726] or [32, Theorem 1]. The proof in [27] relies on the fact that all latin hypercubes of order $v\leq 3$ are linear. Linear hypercubes are easy to count and they are all isotopic. Therefore, the cubes in $\mathcal{C}^{n}(3,2,1)$ are also all isotopic. ∎
Proposition 2.3.
The complement of a $\mathcal{P}$-cube of dimension $n\geq 3$ is not a $\mathcal{P}$-cube.
Proof.
By [22, Proposition 2.3], $\mathcal{P}^{n}(v,k,\lambda)$-cubes have $vk$ incidences ($1$-entries) regardless of the dimension $n$. The number of incidences in the complement is $v^{n}-vk=v(v^{n-1}-k)$. This is too large for any $\mathcal{P}^{n}(v,k^{\prime},\lambda^{\prime})$-cube if $n>2$. ∎
Cubes in $\mathcal{P}^{n}(3,2,1)$ have been classified directly in [22]. The result is reproduced in the first row of Table 1. We see that $\nu(3,2,1)=5$ and the bound (2) is tight in this case. The next feasible parameters are $(7,3,1)$. It is well known that the Fano plane is the unique $(7,3,1)$ design. In [22], two inequivalent $\mathcal{P}^{3}(7,3,1)$-cubes are presented as Examples 2.2 and 2.6. Here we present a full classification of $\mathcal{P}^{n}(7,3,1)$ and $\mathcal{P}^{n}(7,4,2)$-cubes.
Theorem 2.4.
The numbers of cubes in $\mathcal{P}^{n}(7,3,1)$ and $\mathcal{P}^{n}(7,4,2)$ up to equivalence are given in Table 1. In particular, $\nu(7,3,1)=7$ and $\nu(7,4,2)=9$.
$n$
$(v,k,\lambda)$ 2 3 4 5 6 7 8 9 10
$(3,2,1)$ 1 2 1 1 0
$(7,3,1)$ 1 13 20 4 3 2 0 0 0
$(7,4,2)$ 1 877 884 74 19 9 6 5 0
Table 1. Numbers of $\mathcal{P}^{n}(v,k,\lambda)$-cubes up to equivalence.
The bound (2) gives $\nu(7,3,1)\leq 10$ and is not tight for these parameters, but it is tight for $(7,4,2)$. The proof of Theorem 2.4 is a computer calculation that was carried out using the orthogonal array representation of $\mathcal{P}$-cubes. An orthogonal array $OA(N,n,v,t)$ of size $N$, degree $n$, order $v$, and strength $t$ is an $N$-set of $n$-tuples from $\{1,\ldots,v\}^{n}$ such that for any choice of $t$ coordinates, each $t$-tuple from $\{1,\ldots,v\}^{t}$ appears as a restriction of the $n$-tuples to the chosen coordinates exactly $\lambda$ times. This parameter is called the index of the orthogonal array and is given by $\lambda=N/v^{t}$.
A function $C:\{1,\ldots,v\}^{n}\to\{0,1\}$ is the characteristic function of a set of $n$-tuples
$\overline{C}=\{(i_{1},\ldots,i_{n})\in\{1,\ldots,v\}^{n}\mid C(i_{1},\ldots,i_{n})=1\}.$
If $C\in\mathcal{P}^{n}(v,k,\lambda)$, then $\overline{C}$ is an orthogonal array $OA(vk,n,v,1)$ of index $k$ [22, Corollary 2.5]. This means that each element from $\{1,\ldots,v\}$ appears exactly $k$ times in every coordinate. This property is not sufficient; a characterization of $OA(vk,n,v,1)$’s representing $\mathcal{P}^{n}(v,k,\lambda)$-cubes is given in [22, Proposition 2.4]. In the proof of [22, Theorem 2.9] it was also shown that $\overline{C}$ is a distance-invariant non-linear $v$-ary code of length $n$, in which only the distances $\{n-1,n\}$ occur.
The OA-representation of $\mathcal{P}$-cubes is convenient for classification because the size $N=vk$ remains constant across all dimensions $n$. We can take the $21$ incident pairs of the Fano plane and extend them to triples representing $\mathcal{P}^{3}(7,3,1)$-cubes, and continue increasing the dimension by $1$ in this way. If the conditions from [22, Proposition 2.4] are not taken into account, the number of extensions of an $OA(vk,n,v,1)$ is $(vk)!/(k!)^{v}$. For parameters $(7,3,1)$ this is already too large for an exhaustive computer search, even if we do check the conditions for partial extensions using a backtracking algorithm. We amend this by performing isomorph rejection. As an intermediate step, we add the entry $1$ to the new coordinate in (${vk\choose k}$ ways and check the necessary and sufficient conditions. We then eliminate equivalent copies among the valid partial extensions, and proceed by adding the next entry $2$ in the same way. When we reach the last entry $v$, we have all OA-representations of $\mathcal{P}^{n+1}(v,k,\lambda)$-cubes up to equivalence.
We performed this calculation for parameters $(7,3,1)$ and $(7,4,2)$, increasing the dimension until further extension is not possible. We relied on canonical labelings produced by nauty and Traces [26] to eliminate equivalent (partial) OA-representations of cubes. Online versions of Table 1 and other tables in this paper are available on the web page
https://web.math.pmf.unizg.hr/~krcko/results/pcubes.html
The online tables contain links to files with OA-representations of the corresponding cubes that can be used to verify our calculations. The files are given in GAP [12] format. The GAP package Prescribed automorphism groups [20] contains functions for working with $\mathcal{P}$- and $\mathcal{C}$-cubes; see the package manual.
We were not able to perform a full classification for the next parameters $(11,5,2)$, but this might be within reach of today’s computer technology. Instead, we classify $\mathcal{P}^{n}(11,5,2)$-cubes with nontrivial autotopy groups in Theorem 3.8. Computer classification of $\mathcal{C}$-cubes is a much more difficult problem. $\mathcal{C}^{n}(v,k,\lambda)$-cubes can also be represented as orthogonal arrays $OA(kv^{n-1},n,v,n-1)$ of index $k$, see [21, Section 2]. However, the size $N=kv^{n-1}$ grows exponentially with the dimension $n$, so another approach is needed. We have not found a feasible classification strategy even for $\mathcal{C}^{3}(7,3,1)$.
3. Autotopies
Isotopy from an incidence cube to itself is called autotopy. The set of all autotopies of $C$ forms a group with coordinatewise composition, called the full autotopy group and denoted by $\operatorname{Atop}(C)$. This is a generalization of the full automorphism group of a symmetric design. Automorphisms of symmetric designs are usually defined as single permutations of points, because the corresponding permutations of blocks are uniquely determined. This is a consequence of Lemma 2.1: if $A=P_{1}AP_{2}^{t}$ holds for permutation matrices $P_{1}$ and $P_{2}$, then $P_{2}=A^{-1}P_{1}A$. The next two propositions describe how this carries over to higher-dimensional $\mathcal{P}$- and $\mathcal{C}$-cubes.
Proposition 3.1.
Let $(\pi_{1},\ldots,\pi_{n})$ be an autotopy of $C\in\mathcal{P}^{n}(v,k,\lambda)$. Then any component $\pi_{x}$ uniquely determines all other components.
Proof.
For any other component $\pi_{y}$, the pair $(\pi_{x},\pi_{y})$ is an automorphism of the symmetric design $\Pi_{xy}(C)$. Therefore, $\pi_{y}$ is uniquely determined by $\pi_{x}$ and $C$. ∎
Proposition 3.2.
Let $(\pi_{1},\ldots,\pi_{n})$ be an autotopy of $C\in\mathcal{C}^{n}(v,k,\lambda)$. Then any component $\pi_{x}$ is uniquely determined by the $n-1$ other components.
Proof.
Suppose we want to determine $\pi_{n}$ from $\pi_{1},\ldots,\pi_{n-1}$ and $C$. Let $A^{\prime}$ and $A$ be the $2$-sections of $C$ obtained by fixing the first $n-2$ coordinates to $1,\ldots,1$ and $\pi_{1}^{-1}(1),\ldots,\pi_{n-2}^{-1}(1)$, respectively. According to the definition, they are incidence matrices of $(v,k,\lambda)$ designs. The pair $(\pi_{n-1},\pi_{n})$ is an isomorphism between $A$ and $A^{\prime}$:
$A^{\prime}(i,j)$ $=C(1,\ldots,1,i,j)=C(\pi_{1}^{-1}(1),\ldots,\pi_{n-2}^{-1}(1),\pi_{n-1}^{-1}(i),\pi_{n}^{-1}(j))$
$=A(\pi_{n-1}^{-1}(i),\pi_{n}^{-1}(j)),\kern 5.69054pt\forall i,j\in\{1,\ldots,v\}.$
We can write this in matrix form as $A^{\prime}=P_{n-1}AP_{n}^{t}$. Thus, $\pi_{n}$ is uniquely determined by $P_{n}=(A^{\prime})^{-1}P_{n-1}A$. ∎
The previous proposition is best possible, in the sense that $\pi_{x}$ is not determined by fewer than $n-1$ other components of the autotopy. By Theorem 3.4 of [21], a $\mathcal{C}$-cube of dimension $n$ constructed from a $(v,k,\lambda)$ difference set in $G$ has an autotopy group isomorphic to $G^{n-1}$. In this group there are $v$ different choices for $\pi_{n-1}$ and $\pi_{n}$ for any given components $\pi_{1},\ldots,\pi_{n-2}$.
It is known that automorphisms of symmetric $(v,k,\lambda)$ designs fix as many points as blocks. The analogous statement is true for $\mathcal{P}$-cubes.
Proposition 3.3.
Let $\pi\in\operatorname{Atop}(C)$ be an autotopy of $C\in\mathcal{P}^{n}(v,k,\lambda)$. Then every component $\pi_{x}$ has the same number of fixed points.
Proof.
The claim follows from the observation that $(\pi_{x},\pi_{y})$ is an automorphism of the symmetric design $A=\Pi_{xy}(C)$, for all $1\leq x<y\leq n$. If $P_{x}$ and $P_{y}$ are the corresponding permutation matrices, then $A=P_{x}AP_{y}^{t}$ holds. By Lemma 2.1 we can write this as $P_{y}=A^{-1}P_{x}A$. Now $P_{x}$ and $P_{y}$ have the same trace, and this is the number of fixed points of $\pi_{x}$ and $\pi_{y}$. ∎
As a consequence, autotopy groups of $\mathcal{P}$-cubes have the same number of orbits on each coordinate.
Proposition 3.4.
Let $G\leq\operatorname{Atop}(C)$ be an autotopy group of $C\in\mathcal{P}^{n}(v,k,\lambda)$. For any $x\in\{1,\ldots,n\}$, $G_{x}=\{\pi_{x}\mid\pi\in G\}$ is a permutation group acting on $\{1,\ldots,v\}$; the number of orbits of $G_{x}$ does not depend on the choice of $x$.
Proof.
By the Burnside–Cauchy–Frobenius lemma, the number of orbits can be expressed as
$\frac{1}{|G_{x}|}\sum_{\pi_{x}\in G_{x}}f(\pi_{x})=\frac{1}{|G|}\sum_{\pi\in G}f(\pi_{x}).$
Here $f(\pi_{x})$ is the number of fixed points and does not depend on $x$ by Proposition 3.3. ∎
Now it is clear that an autotopy group $G\leq\operatorname{Atop}(C)$ of a $\mathcal{P}$-cube $C$ acts (sharply) transitively on one coordinate if and only if it acts (sharply) transitively on every coordinate; cf. [22, Proposition 3.7]. Furthermore, bounds on the number of fixed points of automorphisms of symmetric $(v,k,\lambda)$ designs apply to all components $\pi_{x}$ of autotopies of $\mathcal{P}^{n}(v,k,\lambda)$-cubes. For example, $f(\pi_{x})\leq v/2$ by [11, Theorem 3] and $f(\pi_{x})\leq k-\sqrt{k-\lambda}$ by [31, Corollary 2.5.6].
On the other hand, autotopies of $\mathcal{C}$-cubes can have components with different numbers of fixed points. The $\mathcal{C}^{3}(7,3,1)$ cube of [21, Example 2.2] has an autotopy of order $7$ with two components fixing no points, and the third component fixing $7$ points. In [21, Propositions 5.1 and 5.3] a total of $2396$ inequivalent $\mathcal{C}^{3}(16,6,2)$-cubes have been constructed. They have autotopies with two components fixing no points, and the third component fixing $f=2$, $4$, $6$, $8$, $12$, $14$, or $16$ points. This also provides counterexamples for bounds on the number of fixed points and other claims stated above.
Our next goal is to classify $\mathcal{P}^{n}(11,5,2)$-cubes with nontrivial autotopy groups. It suffices to consider autotopies of prime orders $p$. By [1, Theorem 2.7], either $p$ divides $v$ or $p\leq k$. The unique $(11,5,2)$ design has full automorphism group of order $660$, hence all possible orders $p\in\{2,3,5,11\}$ occur for dimension $n=2$. Cubes with autotopies of order $11$ are obtained from higher-dimensional $(11,5,2)$ difference sets and have already been classified in [22]; the result is reproduced in the first row of Table 2. Interestingly, each one of these cubes has full autotopy group of order $55$ isomorphic to $\mathbb{Z}_{11}\rtimes\mathbb{Z}_{5}$.
For $p=5$ we adopt the classification strategy from the previous section of successively increasing the dimension. By Proposition 3.3 we know that every component of an autotopy of order $5$ has one fixed point and two cycles of length $5$. Crucially, autotopies are preserved when $\mathcal{P}^{n}(11,5,2)$-cubes are restricted to $n-1$ dimensions by deleting a coordinate in the OA-representation. We can therefore extend the $\mathcal{P}^{n-1}(11,5,2)$-cubes with autotopy of order $5$, respecting the autotopy on the added coordinate and the conditions from [22, Proposition 2.4]. If we do this exhaustively, we get all $\mathcal{P}^{n}(11,5,2)$-cubes with the prescribed autotopy. The procedure was implemented in the C programming language and required a modest amount of CPU time to perform a full classification for $p=5$. Nauty and Traces [26] were used to eliminate equivalent cubes and to compute full autotopy groups.
Proposition 3.5.
Up to equivalence, the numbers of $\mathcal{P}^{n}(11,5,2)$-cubes with an autotopy of order $5$ are given in the second row of Table 2.
$n$
$p$ 2 3 4 5 6 7 8 9 10 11 12
$11$ 2 4 6 6 4 2 1 1 1 0
$5$ 1 283 443 8 7 4 2 1 1 1 0
$3$ 1 4758 0 0 0 0 0 0 0 0 0
$2$ 1 5142 0 0 0 0 0 0 0 0 0
Total 1 10178 443 8 7 4 2 1 1 1 0
Table 2. The $\mathcal{P}^{n}(11,5,2)$-cubes with nontrivial autotopies.
All the new examples from the previous proposition have full autotopy groups of order $5$. Of course, the examples with full autotopy groups of order $55$ were also recovered. The next case $p=3$ was settled by a similar calculation.
Proposition 3.6.
The number of inequivalent $\mathcal{P}^{n}(11,5,2)$-cubes with an autotopy of order $3$ is $4758$ for $n=3$ and 00 for $n\geq 4$; see the third row of Table 2.
Proof.
Automorphisms of order $3$ of the unique $(11,5,2)$ design have $2$ fixed points, each incident with $2$ fixed blocks. By Proposition 3.3 we can assume that all components of the autotopy are $\pi_{x}=(3,4,5)(6,7,8)$ $(9,10,11)$. An exhaustive computer search, systematically extending the $55$ incident pairs of the $(11,5,2)$ design, produced $4758$ inequivalent OA-representations of $\mathcal{P}^{3}(11,5,2)$-cubes invariant under the prescribed autotopy. A few hours of CPU time were required. In the next step none of the $3$-cubes extend to $4$ dimensions. This can be inferred directly by considering the fixed elements. For $n=2$, incidences of the fixed elements are $(1,1)$, $(1,2)$, $(2,1)$, $(2,2)$. The pairs can be extended to triples $(1,1,1)$, $(1,2,2)$, $(2,1,2)$, $(2,2,1)$, but not to quadruples of fixed elements satisfying all requirements. ∎
Five $3$-cubes from the previous proposition have full autotopy groups of order $12$ isomorphic to $(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\rtimes\mathbb{Z}_{3}$. The $4753$ other $3$-cubes have full autotopy groups of order $3$. It remains to classify cubes with autotopies of order $p=2$, i.e. involutions.
Proposition 3.7.
Up to equivalence, the number of $\mathcal{P}^{n}(11,5,2)$-cubes with an involutory autotopy is $5142$ for $n=3$ and 00 for $n\geq 4$; see the fourth row of Table 2.
Proof.
Automorphisms of order $p=2$ of the $(11,5,2)$ design have $3$ fixed points and blocks, forming $3$ incident pairs. Now incidences of the fixed elements can be extended to any dimension. The cubes exist for $n=3$; we classified them by the algorithm described above. The resulting $5142$ inequivalent $3$-cubes cannot be extended to dimension $n=4$. This was established by running the algorithm in parallel and required some $2$ years of CPU time in total. Not a single extension was found, so $\mathcal{P}^{n}(11,5,2)$-cubes with an involutory autotopy do not exist for $n\geq 4$. ∎
The five $\mathcal{P}^{3}(11,5,2)$-cubes with full autotopy groups of order $12$ were also found in Proposition 3.7. Of the remaining cubes, $71$ have full autotopy groups of order $4$ isomorphic to $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and $5066$ have full autotopy groups of order $2$. We calculated the total numbers of $\mathcal{P}^{n}(11,5,2)$-cubes with nontrivial autotopies by concatenating the lists for $p=11$, $5$, $3$, $2$ and eliminating equivalent copies.
Theorem 3.8.
Cubes in $\mathcal{P}^{n}(11,5,2)$ with nontrivial autotopy groups exist if and only if $2\leq n\leq 11$. The number of such cubes up to equivalence is $1$, $10178$, $443$, $8$, $7$, $4$, $2$, $1$, $1$, and $1$ for successive dimensions $n$ in this range.
In [21] and [22], constructions of $C^{3}(16,6,2)$ and $\mathcal{P}^{3}(16,6,2)$-cubes with prescribed autotopy groups $G$ were performed by a different approach. Essentially, all orthogonal arrays $OA(1536,3,16,2)$ and $OA(96,$ 3,16,1)3,16,1)3 , 16 , 1 ) invariant under $G$ were constructed and the ones not representing $\mathcal{C}$- and $\mathcal{P}$-cubes were discarded. Because of high proportions of “bad” OAs, rather large prescribed groups had to be used: $|G|\geq 512$ in [21] and $|G|\geq 18$ in [22]. We now have a more efficient approach for $\mathcal{P}$-cubes and can handle smaller groups. As an example, we present the following result.
Proposition 3.9.
There are exactly $1076$ inequivalent $\mathcal{P}^{3}(16,6,2)$-cubes with an autotopy of order $8$ acting in two cycles on each coordinate.
$(T,P)$
 $(\mathcal{D}_{R},\mathcal{D}_{R},\mathcal{D}_{R})$
 $(\mathcal{D}_{R},\mathcal{D}_{R},\mathcal{D}_{G})$
 $(\mathcal{D}_{R},\mathcal{D}_{R},\mathcal{D}_{B})$
 $(\mathcal{D}_{R},\mathcal{D}_{G},\mathcal{D}_{G})$
 $(\mathcal{D}_{R},\mathcal{D}_{G},\mathcal{D}_{B})$
 $(\mathcal{D}_{R},\mathcal{D}_{B},\mathcal{D}_{B})$
 $(\mathcal{D}_{G},\mathcal{D}_{G},\mathcal{D}_{G})$
 $(\mathcal{D}_{G},\mathcal{D}_{G},\mathcal{D}_{B})$
 $(\mathcal{D}_{G},\mathcal{D}_{B},\mathcal{D}_{B})$
 $(\mathcal{D}_{B},\mathcal{D}_{B},\mathcal{D}_{B})$
Total
$(8,1)$ 4 48 76 124 152 56 102 136 48 35 781
$(8,2)$ 21 0 8 72 0 64 0 8 0 6 179
$(8,3)$ 0 0 0 0 0 0 6 0 0 5 11
$(8,6)$ 1 0 0 0 0 0 0 0 0 2 3
$(16,1)$ 23 8 0 16 0 0 12 0 0 0 59
$(16,2)$ 18 0 0 8 0 0 0 0 0 0 26
$(16,3)$ 0 0 0 0 0 0 4 0 0 0 4
$(16,6)$ 4 0 0 0 0 0 0 0 0 0 4
$(32,1)$ 1 0 0 0 0 0 0 0 0 0 1
$(32,2)$ 3 0 0 0 0 0 0 0 0 0 3
$(32,3)$ 1 0 0 0 0 0 0 0 0 0 1
$(32,6)$ 1 0 0 0 0 0 0 0 0 0 1
$(48,2)$ 1 0 0 0 0 0 0 0 0 0 1
$(48,6)$ 1 0 0 0 0 0 0 0 0 0 1
$(96,6)$ 1 0 0 0 0 0 0 0 0 0 1
Total 80 56 84 220 152 120 124 144 48 48 1076
Table 3. $\mathcal{P}^{3}(16,6,2)$-cubes with an autotopy of order $8$ acting semiregularly. Table of the statistics of the $1076$ cubes in Proposition 3.9.
Proof.
Three $(16,6,2)$ designs exist [24]. We will call them the red, green and blue design and denote them by $\mathcal{D}_{R}$, $\mathcal{D}_{G}$, and $\mathcal{D}_{B}$. The full automorphism groups are of orders $|\operatorname{Aut}(\mathcal{D}_{R})|=11520$, $|\operatorname{Aut}(\mathcal{D}_{G})|=768$, and $|\operatorname{Aut}(\mathcal{D}_{B})|=384$. All three designs have automorphisms of order $8$ acting in two cycles of length $8$ on the points and blocks. We ran the dimension increasing algorithm and found that they extend to $440$, $744$, and $596$ inequivalent $\mathcal{P}^{3}(16,6,2)$-cubes invariant under the prescribed autotopy, respectively. The total number of extensions was determined by concatenating the lists and eliminating equivalent cubes. ∎
Figure 3. A $\mathcal{P}^{3}(16,6,2)$-cube with non-isomorphic projections.
A detailed statistics of the $1076$ cubes is given in Table 3. The cubes are divided according to the size of the full auto(para)topy group and the $2$-projections. The full autoparatopy group $\operatorname{Apar}(C)$ contains all combinations of isotopy and conjugation mapping $C$ onto itself. The group size is given as $(T,P)$, where $T=|\operatorname{Atop}(C)|$ and $P=|\operatorname{Apar}(C)|/T$. A question whether $\mathcal{P}^{3}(16,6,2)$-cubes with three non-isomorphic projections exist was raised in [22]. From the table, we see that there are $152$ cubes with projections $(\mathcal{D}_{R},\mathcal{D}_{G},\mathcal{D}_{B})$ and the prescribed autotopy of order $8$. One example is shown in Figure 3, with red, green, and blue light sources so that the projections are in the appropriate color.
4. Difference sets
Let $G$ be an additively written group of order $v$, not necessarily abelian. Henceforth we will index cubes with the elements of $G$ instead of the integers $\{1,\ldots,v\}$. Thus, $\mathcal{C}$- and $\mathcal{P}$-cubes are functions $C:G^{n}\to\{0,1\}$ with all $2$-sections or $2$-projections being incidence matrices of symmetric $(v,k,\lambda)$ designs.
A $(v,k,\lambda)$ difference set in $G$ is a $k$-subset $D\subseteq G$ such that every element $g\in G\setminus\{0\}$ can be written as $g=a-b$ with $a,b\in D$ in exactly $\lambda$ ways. By [21, Theorem 3.1], difference sets give rise to $\mathcal{C}$-cubes of arbitrary dimension $n$. Using the Iverson bracket, a cube $C\in\mathcal{C}^{n}(v,k,\lambda)$ is given by
$C(g_{1},\ldots,g_{n})=[g_{1}+\ldots+g_{n}\in D].$
This is a special case of [5, Theorem 2.9]. Properties of these difference cubes were studied in Section 3 of [21]. In Section 4, the construction was generalized to the so-called group cubes [21, Theorem 4.1]. This construction gives examples not equivalent to any difference cube, having non-isomorphic $(v,k,\lambda)$ designs as $2$-sections. The construction of [21, Theorem 4.1] still requires difference sets; the only known examples of $\mathcal{C}$-cubes not coming from this constructions are the $\mathcal{C}^{3}(16,6,2)$-cubes of [21, Proposition 5.3].
A stronger kind of difference sets are needed for $\mathcal{P}$-cubes. An $n$-dimensional $(v,k,\lambda)$ difference set [22, Definition 3.1] is a $k$-subset of $n$-tuples $D\subseteq G^{n}$ such that $\{d_{x}-d_{y}\mid d\in D\}$ is an “ordinary” $(v,k,\lambda)$ difference set for every pair of coordinates $1\leq x<y\leq n$. By [22, Proposition 3.2], the development
$\operatorname{dev}D=\{(d_{1}+g,\ldots,d_{n}+g)\mid g\in G,\,d\in D\}\subseteq G^{n}$
is an OA-representation $\overline{C}$ of a cube $C\in\mathcal{P}^{n}(v,k,\lambda)$ indexed by $G$. Properties of these cubes were studied in Section 3 of [22]. There, it was shown that $n$-dimensional difference sets can be normalized so that all $n$-tuples in $D$ start with a 00 coordinate. Using a different kind of normalization, we will now prove a bound on the dimension of $D$ stronger than the bound (2).
Theorem 4.1.
If an $n$-dimensional $(v,k,\lambda)$ difference set $D\subseteq G^{n}$ exists, then $n\leq v$.
Proof.
For a group element $g\in G$ and an $n$-tuple $d=(d_{1},\ldots,d_{n})\in G^{n}$, denote by $g+_{x}d=(d_{1},\ldots,d_{x-1},g+d_{x},d_{x+1},\ldots,d_{n})$. We claim that $D^{\prime}=\{g+_{x}d\mid d\in D\}$ is also an $n$-dimensional $(v,k,\lambda)$ difference set. For any other index $y\in\{1,\ldots,n\}$, the set $\{d_{x}-d_{y}\mid d\in D^{\prime}\}=\{g+d_{x}-d_{y}\mid d\in D\}$ is a left translate of the difference set $\{d_{x}-d_{y}\mid d\in D\}$, hence also a $(v,k,\lambda)$ difference set. By repeated application of this transformation, we can make an $n$-dimensional difference set $D^{\prime}$ containing the $n$-tuple $(0,\ldots,0)$. Now any other $n$-tuple $d\in D^{\prime}$ must have distinct coordinates, because $d_{x}=d_{y}$ would imply that the set of differences $\{d_{x}-d_{y}\mid d\in D^{\prime}\}$ contains fewer than $k$ elements of $G$. Therefore, the number of coordinates $n$ cannot exceed $v=|G|$. ∎
The normalization of $n$-dimensional difference sets from [22] does not change the development $\operatorname{dev}D$. The new normalization can change the development, but $\operatorname{dev}D$ and $\operatorname{dev}D^{\prime}$ are always isotopic. In [22, Propositions 3.5 and 3.7], $\mathcal{P}$-cubes coming from difference sets were characterized as having an autotopy group acting regularly on the coordinates. Theorem 2.4 shows that the bound $n\leq v$ does not hold for general $\mathcal{P}^{n}(v,k,\lambda)$-cubes.
The largest integer $n$ such that an $n$-dimensional $(v,k,\lambda)$ difference set in $G$ exists was denoted by $\mu_{G}(v,k,\lambda)$ in [22]. To determine values of this function, a computer classification of small $n$-dimensional difference sets was performed using the library of difference sets available in the GAP package DifSets [29]. In [22], calculations were performed in GAP. Exact values of $\mu_{G}(16,6,2)$ were determined in [22, Table 3] for $10$ of the $14$ groups of order $16$, and lower bounds were given for the remaining $4$ groups.
We have implemented the classification algorithm in the C programming language and determined more exact values of $\mu$ and better lower bounds. Detailed results of our calculations are presented in Table 4. The newly established values of $\mu$ are summarized in Theorem 4.2. When there is only one group of order $v$ up to isomorphism we write $\mu(v,k,\lambda)$, and when there are several groups we identify them by their ID in the GAP library of small groups [12].
Theorem 4.2.
For groups of order $16$ with IDs $10$ and $13$ the values of $\mu_{G}(16,6,2)$ are $4$ and $8$. For groups with IDs $2$, $3$, $4$, $5$, $6$, and $8$ the values of $\mu_{G}(16,10,6)$ are $6$, $6$, $6$, $5$, $4$, and $6$, respectively. Furthermore, $\mu(13,9,6)=13$, $\mu(19,9,4)=\mu(19,10,5)=19$, $\mu(23,11,5)=\mu(23,12,6)=23$, and $\mu(31,6,1)=31$.
$n$
$G$ $(v,k,\lambda)$ 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19
$\mathbb{Z}_{7}$ $(7,3,1)$ 1 2 2 1 1 1
$(7,4,2)$ 1 2 2 1 1 1
$\mathbb{Z}_{11}$ $(11,5,2)$ 1 2 4 6 6 4 2 1 1 1
$(11,6,3)$ 1 2 4 6 6 4 2 1 1 1
$\mathbb{Z}_{13}$ $(13,4,1)$ 1 3 7 10 14 14 10 7 3 1 1 1
$(13,9,6)$ 1 146 422 652 305 60 13 8 3 1 1 1
$\mathbb{Z}_{15}$ $(15,7,3)$ 1 3 0 0 0 0 0 0 0 0 0 0 0 0
$(15,8,4)$ 1 6 1 0 0 0 0 0 0 0 0 0 0 0
ID2 $(16,6,2)$ 1 31 81 0 0 0 0 0 0 0 0 0 0 0 0
$(16,10,6)$ 1 2565 152314 12115 36 0 0 0 0 0 0 0 0 0 0
ID3 $(16,6,2)$ 1 16 55 0 0 0 0 0 0 0 0 0 0 0 0
$(16,10,6)$ 1 6638 462880 111294 196 0 0 0 0 0 0 0 0 0 0
ID4 $(16,6,2)$ 1 38 113 0 0 0 0 0 0 0 0 0 0 0 0
$(16,10,6)$ 1 6516 389060 34076 53 0 0 0 0 0 0 0 0 0 0
Table 4. Numbers of inequivalent $\mathcal{P}^{n}(v,k,\lambda)$-cubes obtained from difference sets.
$n$
$G$ $(v,k,\lambda)$ 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19
ID5 $(16,6,2)$ 2 56 140 0 0 0 0 0 0 0 0 0 0 0 0
$(16,10,6)$ 2 10680 323520 6874 0 0 0 0 0 0 0 0 0 0 0
ID6 $(16,6,2)$ 1 8 6 0 0 0 0 0 0 0 0 0 0 0 0
$(16,10,6)$ 1 506 1192 0 0 0 0 0 0 0 0 0 0 0 0
ID8 $(16,6,2)$ 1 18 44 0 0 0 0 0 0 0 0 0 0 0 0
$(16,10,6)$ 1 3746 76580 5444 8 0 0 0 0 0 0 0 0 0 0
ID9 $(16,6,2)$ 1 38 112 0 0 0 0 0 0 0 0 0 0 0 0
ID10 $(16,6,2)$ 1 86 1941 0 0 0 0 0 0 0 0 0 0 0 0
ID11 $(16,6,2)$ 1 24 88 0 0 0 0 0 0 0 0 0 0 0 0
ID13 $(16,6,2)$ 2 129 4960 19734 8106 374 2 0 0 0 0 0 0 0 0
$\mathbb{Z}_{19}$ $(19,9,4)$ 1 8 14 36 86 154 228 280 280 228 154 86 36 14 4 1 1 1
$(19,10,5)$ 1 8 14 36 86 154 228 280 280 228 154 86 36 14 4 1 1 1
$\mathbb{Z}_{21}$ $(21,5,1)$ 1 2 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 $\cdots$
$F_{21}$ $(21,5,1)$ 1 6 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 $\cdots$
Table 4. Numbers of inequivalent $\mathcal{P}^{n}(v,k,\lambda)$-cubes obtained from difference sets (continued).
$n$
$G$ $(v,k,\lambda)$ 2 3 4 5 6 7 8 9 10 11
$\mathbb{Z}_{23}$ $(23,11,5)$ 1 11 20 69 207 492 984 1630 2282 2694
$(23,12,6)$ 1 11 20 69 207 492 984 1630 2282 2694
$\mathbb{Z}_{31}$ $(31,6,1)$ 1 10 49 195 812 2846 8528 21731 47801 91148
$n$
$G$ $(v,k,\lambda)$ 12 13 14 15 16 17 18 19
$\mathbb{Z}_{23}$ $(23,11,5)$ 2694 2282 1630 984 492 207 69 20
$(23,12,6)$ 2694 2282 1630 984 492 207 69 20
$\mathbb{Z}_{31}$ $(31,6,1)$ 151924 221959 285357 323396 323396 285357 221959 151924
$n$
$G$ $(v,k,\lambda)$ 20 21 22 23 24 25 26 27 28 29 30 31
$\mathbb{Z}_{23}$ $(23,11,5)$ 4 1 1 1
$(23,12,6)$ 4 1 1 1
$\mathbb{Z}_{31}$ $(31,6,1)$ 91148 47801 21731 8528 2846 811 187 38 6 1 1 1
Table 4. Numbers of inequivalent $\mathcal{P}^{n}(v,k,\lambda)$-cubes obtained from difference sets (continued).
The previously known values of $\mu$ are given in [22, Tables 2 and 3] and can also be read from Table 4. The groups of order $16$ with IDs $1$, $7$, $12$, and $14$ are missing from Table 4. The former two groups are the cyclic group $\mathbb{Z}_{16}$ and the dihedral group $D_{16}$, which do not contain difference sets. For the latter two groups we could not completely classify $n$-dimensional difference sets due to large numbers of inequivalent $\mathcal{P}$-cubes arising from them. Example 4.3 establishes an improved lower bound $\mu_{G}(16,6,2)\geq 11$ for the group with ID $12$. The group with ID $14$ is the elementary abelian group $\mathbb{Z}_{2}^{4}$, and will be dealt with in Corollary 4.5.
Example 4.3.
The group of order $16$ with GAP ID $12$ is isomorphic to $G=\mathbb{Z}_{2}\times Q_{8}$, where $Q_{8}$ is the quaternion group. If elements of the factor groups are $\mathbb{Z}_{2}=\{0,1\}$ and $Q_{8}=\{1,i,j,k,-1,-i,-j,-k\}$, the following is a $11$-dimensional $(16,6,2)$ difference set in $G$:
[…]
$((0,1),(0,j),(1,1),(0,1),(0,-i),(1,1),(0,j),(0,k),(1,1),(1,i),(1,-j)),$
$((0,1),(0,i),(0,j),(1,1),(1,-j),(0,i),(1,1),(0,j),(0,-1),(1,1),(1,i)),$
$((0,1),(1,1),(0,i),(1,-k),(0,1),(0,-i),(0,-1),(0,-1),(0,j),(1,-i),(0,j)),$
$((0,1),(0,-1),(1,-k),(0,-1),(1,1),(0,1),(0,1),(1,i),(0,k),(1,-j),(0,1)),$
$((0,1),(1,-k),(0,-1),(0,i),(0,j),(1,-j),(1,i),(0,1),(0,1),(0,j),(1,1))\,\}.$
We see that equality is reached quite often in the upper bound $\mu_{G}(v,k,\lambda)\leq v$ of Theorem 4.1. For parameters $(q,(q-1)/2,(q-3)/4)$, this is explained by the construction of $q$-dimensional Paley difference sets in the additive group of $\mathbb{F}_{q}$, $q\equiv 3\pmod{4}$; see [22, Theorem 4.1]. It was noted that the same construction works for cyclotomic difference sets ($4$th and $8$th powers in $\mathbb{F}_{q}$ for appropriate orders $q$), but this does not explain the equality $\mu(13,4,1)=\mu(13,9,6)=13$. We will now generalize this construction to difference sets with arbitrary parameters in elementary abelian groups. The key observation in the proof of the following theorem is that group automorphisms $\varphi:G\to G$ preserve difference sets in $G$.
Theorem 4.4.
Let $G$ be an elementary abelian group, i.e. the additive group of a finite field $\mathbb{F}_{q}$. Then any $(q,k,\lambda)$ difference set in $G$ extends to dimension $q$.
Proof.
Let $D=\{d_{1},\ldots,d_{k}\}\subseteq G$ be a $(q,k,\lambda)$ difference set, $\alpha\in\mathbb{F}_{q}$ a primitive element, and $w=(0,1,\alpha,\alpha^{2},\ldots,\alpha^{q-2})\in G^{q}$. An extension of $D$ to dimension $q$ is $\vec{D}=\{d_{1}w,\ldots,d_{k}w\}$, where each coordinate of $w$ is multiplied by an element of $D$. To prove that $\vec{D}$ is indeed a $q$-dimensional $(q,k,\lambda)$ difference set, we have to check that the differences of any two coordinates are difference sets in $G$. The first coordinate of any vector in $\vec{D}$ is 00, and if we subtract the coordinate with $\alpha^{y}$ in $w$, we get $\{-\alpha^{y}d_{1},\ldots,-\alpha^{y}d_{k}\}$. This is a $(q,k,\lambda)$ difference set in $G$ because it is the image of $D$ by the group automorphism $\varphi:G\to G$, $\varphi(g)=-\alpha^{y}g$. Next, we subtract two non-zero coordinates, say with $\alpha^{x}$ and $\alpha^{y}$ in $w$ for some $x<y$. We get the set
$\{\alpha^{x}(1-\alpha^{y-x})d_{1},\ldots,\alpha^{x}(1-\alpha^{y-x})d_{k}\}.$
This is again the image of $D$ by a group automorphism $\psi:G\to G$, $\psi(g)=\alpha^{x}(1-\alpha^{y-x})g$. ∎
Two missing values of $\mu_{G}$ follow from Theorem 4.4.
Corollary 4.5.
For $G=\mathbb{Z}_{2}^{4}$, $\mu_{G}(16,6,2)=\mu_{G}(16,10,6)=16$.
A further generalization of Theorem 4.4 applies to arbitrary groups $G$. An antiautomorphism of $G$ is a bijection $\varphi:G\to G$ such that
$\varphi(g+h)=\varphi(h)+\varphi(g),\kern 8.53581pt\forall g,h\in G.$
If $G$ is abelian, automorphisms and antiautomorphisms coincide. Antiautomorphisms of non-abelian groups are the functions $-\varphi$, where $\varphi$ is an automorphism. Thus, antiautomorphisms of $G$ also map $(v,k,\lambda)$ difference set $D\subseteq G$ to difference sets with the same parameters.
Definition 4.6.
We say that $R=\{\varphi_{1},\ldots,\varphi_{n-1}\}$ is a regular set of (anti)automorphisms of $G$ if each $\varphi_{x}:G\to G$ is an automorphism or antiautomorphism, and each difference $\varphi_{x}-\varphi_{y}$ is an automorphism or antiautomorphism for $11\leq x<y\leq n-$ 1.
Given such a set $R$ of size $n-1$ and an element $d\in D$, denote by $\vec{d}$ the $n$-tuple $(0,\varphi_{1}(d),\ldots,\varphi_{n-1}(d))\in G^{n}$. Then the set $\vec{D}=\{\vec{d}\mid d\in D\}$ is an $n$-dimensional $(v,k,\lambda)$ difference set. This proves the following theorem.
Theorem 4.7.
If a group $G$ allows a regular set of (anti)automorphisms of size $n-1$, then any $(v,k,\lambda)$ difference set in $G$ extends to dimension $n$.
Theorem 4.4 can be seen as a special case of Theorem 4.7 by using multiplication in $\mathbb{F}_{q}$. The functions $\varphi_{x}:\mathbb{F}_{q}\to\mathbb{F}_{q}$, $\varphi_{x}(g)=\alpha^{x}g$ for $x=0,\ldots,q-2$ constitute a regular set of automorphisms of the additive group, yielding the construction of Theorem 4.4. Kerdock sets provide more examples of regular sets in elementary abelian $2$-groups [3, 7, 15, 16, 18] and $3$-groups [28].
Theorem 4.7 also applies to groups that are not elementary abelian. For example, $\varphi_{1}(g)=g$ and $\varphi_{2}(g)=2g$ are automorphisms of the group $\mathbb{Z}_{15}$. The difference $(\varphi_{1}-\varphi_{2})(g)=-g$ is an automorphism as well. Thus, $\{\varphi_{1},\varphi_{2}\}$ is a regular set and all $(15,7,3)$ and $(15,8,4)$ difference sets extend at least to dimension $3$. In Table 4, we see that there is a $4$-dimensional $(15,8,4)$ difference set, but it cannot be obtained from Theorem 4.7 because $\mathbb{Z}_{15}$ does not allow regular sets of size $3$. The next result applies to difference sets in arbitrary cyclic groups $\mathbb{Z}_{v}$.
Corollary 4.8.
Let $p$ be the smallest prime divisor of $v$. Then any cyclic $(v,k,\lambda)$ difference set extends at least to dimension $p$.
Proof.
Automorphisms of $\mathbb{Z}_{v}$ are of the form $\varphi_{a}(g)=ag$, where $a\in\mathbb{Z}_{v}\setminus\{0\}$ is relatively prime to $v$. The set $\{\varphi_{1},\ldots,\varphi_{p-1}\}$ is a regular set of size $p-1$. Hence, any difference set in $\mathbb{Z}_{v}$ extends to dimension $p$ by Theorem 4.7. ∎
For a non-abelian example, we turn to groups of order $27$. There are five such groups, but only two of them contain nontrivial difference sets [29]. These are the elementary abelian group $\mathbb{Z}_{3}\times\mathbb{Z}_{3}\times\mathbb{Z}_{3}$ (GAP group ID 5) and a semidirect product $\mathbb{Z}_{9}\rtimes\mathbb{Z}_{3}$ (GAP group ID 4). Difference sets in the former group extend to dimension $27$ by Theorem 4.4. A regular set in the latter group is given in the next example, using multiplicative notation.
Example 4.9.
The group of order $27$ with GAP ID $4$ can be presented as $G=\langle a,b\mid a^{9}=b^{3}=1,\,ba=a^{4}b\rangle$. Let $\varphi_{1}(g)=g$ be the identity automorphism of $G$ and $\varphi_{2}(g)=g^{-1}$ the corresponding antiautomorphism. The “difference” $\varphi_{1}(g)(\varphi_{2}(g))^{-1}=g^{2}$ is an antiautomorphism, namely the one corresponding to $\varphi\in\operatorname{Aut}(G)$, $\varphi(a)=a^{7}$, $\varphi(b)=b$. Thus, $\{\varphi_{1},\varphi_{2}\}$ is a regular set of (anti)automorphisms of $G$.
By Theorem 4.7, all $(27,13,6)$ and $(27,14,7)$ difference sets in $G=\langle a,b\mid a^{9}=b^{3}=1,\,ba=a^{4}b\rangle$ extend to dimension $3$.
5. Final observations
A curious fact visible in Table 4 is equality of the numbers for some complementary parameters. Ordinary ($2$-dimensional) difference sets and designs are rarely discussed for both sets of complementary parameters, because complementation is a bijection. This is also true for higher-dimensional $\mathcal{C}$-cubes, but not for $\mathcal{P}$-cubes as we have seen in Proposition 2.3. However, the numbers in Table 4 coincide for complementary Paley-Hadamard parameters $\mathcal{P}^{n}(4m-1,2m-1,m-1)$ and $\mathcal{P}^{n}(4m-1,2m,m)$, when $4m-1=p^{s}$ is a prime power and $G=(\mathbb{Z}_{p})^{s}$ is an elementary abelian group. For small parameters, this can be explained by the following bijection.
Every $(4m-1,2m-1,m-1)$ difference set $D\subseteq G$ can be uniquely extended to a $(4m-1,2m,m)$ difference set $D\cup\{a\}$ by adding the element $a=-2\sum_{d\in D}d$. We have checked that this is a bijection between $(4m-1,2m-1,m-1)$ and $(4m-1,2m,m)$ difference sets in the groups $\mathbb{Z}_{7}$, $\mathbb{Z}_{11}$, $\mathbb{Z}_{19}$, and $\mathbb{Z}_{23}$ from Table 4, and also in $G=(\mathbb{Z}_{3})^{3}$. We don’t have a proof for arbitrary elementary abelian groups. However, if this is indeed a bijection between ordinary difference sets, then it is easy to establish a $1$-to-$1$ correspondence between $n$-dimensional difference sets. Given an $n$-dimensional $(4m-1,2m-1,m-1)$ difference set $D\subseteq G^{n}$, one simply adds the $n$-tuple $\left(-2\sum_{d\in D}d_{1},\ldots,-2\sum_{d\in D}d_{n}\right)$ to obtain an $n$-dimensional $(4m-1,2m,m)$ difference set.
If a $(4m-1,2m-1,m-1)$ difference set $D\subseteq G$ extends to a $(4m-1,2m,m)$ difference set $D\cup\{a\}$, then the complement $D^{\prime}=G\setminus(D\cup\{a\})$ is also a $(4m-1,2m-1,m-1)$ difference set in $G$. Then $(D-a)\cup(D^{\prime}-a)\cup\{0\}$ is a tiling of $G$ by two difference sets, and $D-a$ is a skew Hadamard difference set by [4, Theorem 8]. A long-standing conjecture was that the classical Paley difference sets are the only skew Hadamard difference sets in elementary abelian groups, but it has been refuted in [8, 9] for groups of orders $3^{m}$, $m\geq 5$ odd. It might also be the case that our bijection holds only for elementary abelian groups of small orders, but we don’t have a counterexample.
Yet another apparent symmetry in Table 4 is equality of the numbers for “complementary dimensions” $n$ and $v-n$. Equality holds for some small parameters $(v,k,\lambda)$ when the upper bound of Theorem 4.1 is reached: $(7,3,1)$, $(7,4,2)$, $(11,5,2)$, $(11,6,3)$, and $(13,4,1)$. It breaks for the complementary parameters $(13,9,6)$, while for $(19,9,4)$, $(19,10,5)$, $(23,11,5)$, and $(23,12,6)$ equality holds except for dimension $n=3$. For $(31,6,1)$, it breaks for $n=3$, $4$, $5$, $6$ and holds for the remaining dimensions. We don’t have an explanation for this phenomenon.